For a lead�tin alloy of composition 80 wt% Sn�20 wt% Pb and at 180�C (355�F) do the following: (a) Determine the mass fractions of the ? and ? phases. (b) Determine the mass fractions of primary ? and eutectic microconstituents. (c) Determine the mass fraction of eutectic ?.
Answer #1

Composition (at% Sn) 0 20 40 60 100 327 C 600 300 Liquid 500 232 C 200 183°C 8.3 61.9 97.8 100 ?+? 100 2 0 0 20 40 60 100 (Pb) Composition (wt% Sn) (Sn)

a). W\alpha = 97.8 - 80 / 97.8 - 18.3

= 0.224

W\beta = 1 -  W\alpha

= 1 - 0.224

= 0.776

b). mass fraction of primary, W\beta' = 80 - 61.9 / 97.8 - 61.9

= 0.504

mass fraction of eutectic, We = 1 - W\beta'

= 1 - 0.504

= 0.496